OpenStudy (anonymous):

how many ounces of pure sugar must be added to 80 ounces of a 28% sugar solution to obtain a 36% solution

OpenStudy (anonymous):

what is 28% of 80?

OpenStudy (anonymous):

22.4

OpenStudy (anonymous):

ok good, so at the moment you have \(22.4\) ounces of sugar in your \(80\) ounces of mix add \(x\) ounces and you will have \(22.4+x\) ounces of sugar, and \(80+x\) ounces of mix you want the \(80+x\) ounces to be \(36\%\) sugar, set \[22.4+x=.36(80+x)\] and solve for \(x\)

OpenStudy (anonymous):

may be easier to solve if you get rid of the annoying decimals by multiplying both sides by 100 and start with \[2240+100x=36(80+x)\]